OpenStudy (anonymous):

solve for x.squareroot 7-3x =1

OpenStudy (anonymous):



OpenStudy (anonymous):

7-3x = 1^2 7-3x = 1 7-1 = 3x 6 = 3x x = 6/3 x = 2

OpenStudy (anonymous):

thank you :)

OpenStudy (australopithecus):

\[\sqrt{7-3x} = 1\] is the same as (7-3x)^(1/2) = 1 raise both sides to the power of (7-3x)^((1/2)(2)) = 1^(2) 7 - 3x = 1 - 3x = 7 x = -7/3

OpenStudy (australopithecus):

now pet your invisible cat